UT Nov 18 2021 - D/E Night



Night Info

  • Sunset 17:22
  • Twilight Ends 18:43
  • Twilight Starts 05:27
  • Moon sets 06:17
  • Sunrise 06:47

Night Plan

  • Starting Config: LBTI
  • DONE LBTI Pallas Observations -3hrs
    • 18:15-21:15hr
    • 30deg elevation limit 22:12hr local time
    • attempt up to 1mag clouds, 1.5" seeing
  • Reconfigure to MODS-bino - 21:15-21:30hr
    • DONE M31 Nova 21:30-22:45hr
    • any(ish) conditions
  • Reconfigure for LUCI-bino 22:45-23hr
    • LUCI eng testing -2hr -23hr-01hr
      • DONE LUCI Sub-frame readout tests (bright object out of readout)
      • DONE ADI tests
      • DONE LUCI stress tests RTD/LUCI software updates
    • DONE Arp 187 - 1hr -01-02hr
      • hits 30deg elevation 04hr
  • Reconfigure to LBCs 02-02:30hr
    • LBC timing tests


Staff

  • OSA: J.Williams
  • ISp: E. Solheid
  • Mtn Manager: M. Wagner
  • ISA: B.Rothberg, R.T. Gatto

Night Log

Closed Dome

  • UT 22:53 (-1 day) - waking MODS and simSnap
    • OK
  • UT 22:59 (-1 day) - starting LUCI GUIs, test dark, init all
    • L1 0011-0012 - 2.51sec dark
    • L2 0001-0002 - 2.51sec dark
    • putting N30fieldstop away
    • Checking FieldStops for both LUCIs
    • L1 0013-0014 - OK
    • L2 0003-0004 - OK
  • UT 23:03 - starting up LBCs - test biases
    • OK
  • UT 00:20 - opening dome
  • UT 02:26 - shutters open ready to go


LBTI - Pallas Observations

  • UT 00:36 - sky too bright for Pallas, running tests first
  • UT 00:59 - loops close, NCPA, GOpt - started science collection
  • UT 01:25 - taking PSF star data
  • UT 01:34 - returning to Pallas
  • will continue to observe until close to 30deg elevation limit
  • UT 04:21 - safe skiping
  • UT 04:24 - AO loops closed again
  • UT 04:33 - finished LMIRCam data
  • next to test ALES observations until elevation limit
  • skipping GOpt and NCAP for ALES test
  • UT 04:56 - finished, reconfiguring for MODS

MODS - M31 Nova RMWagner

  • putting in Observing mode, homing AGWs
  • go to target first, then do G191-B2B
  • UT 05:10 - pointing check
  • UT 05:17 - sending collimation preset
  • UT 05:20 - sending preset - using 1" wide slit (seeing on collimation was ~0.9")
    • acquisition field image is 120sec (vs 60sec previously) in case nova has dimmed significantly
    • clouds unexpectedly thickening up
    • M1 offsetxy -0.317 10.572 rel
    • M2 offsetxy 4.223 8.595 rel
    • 120sec helped with MODS1, MODS2 was bright enough to centroid on
    • M1 OK
    • M2 OK
  • UT 05:32 - starting spectroscopy
  • UT 05:34 - Guiders 1.1", 1.3" (and here I was thinking I was TOO conservative using the 1" slit)
  • UT 05:43 - seeing in guiders back down to <1" - must be the calming effect of the clouds
  • UT 05:56 - can see the H-alpha in both MODS in red channel, also can see the continuum trace in the blue channels, both MODS
  • UT 05:57 - clouds continue to thcken, seeing is getting softer 1.35"-1.5" on guiders
  • UT 05:59 - guiders showing 2"+ seeing! uhoh!
  • UT 06:03 - seeing is variable
  • UT 06:19 - 2nd exposure still see h-alpha and multiple emissions lines and continuum in red channels - blue is iffy...
  • UT 06:31 - seeing did get better, bounced between 2" and 1.3" - getting 0.5-1mag dips in flux from clouds passing over
  • UT 06:46 - seeing improved - back to ~1"
  • UT 06:51 - or about 1.1"-1.2" - still having 0.5-1mag dips in flux
  • UT 07:00 - taking 1 more exposure...conditions are now significantly better (seeing-wise)
    • M1 had a slow readout, it will lag behind the rest by about 1min
    • had to run blue expdone on MODS-2 - didn't spit out the last file
    • M1B 0003-0007
    • M1R 0006-0010
    • M2B 0003-0007
    • M2R 0006-0010
    • just noticed odd error - however, the files are in /newdata and no error appeared in the GUI or popups - this was seen in the mods2Disp terminal
    • * ERROR: Cannot read FITS header of
      Reason: IOError(21, 'Is a directory')
  • UT 07:26 - sending preset to G191B2B
    • M1 offsetxy -0.276 11.022 rel
    • M2 offsetxy 4.696 8.639 rel
    • M1 looks good
    • M2 tweak offsetxy -0.958 -0.327 rel
  • UT 07:35 - starting spectra
  • UT 07:49 Reconfigure for bino-LUCIs, putting MODS to sleep done for the night with them

LUCI-Eng Tests

ADI Test
  • FW Tau
  • UT 08:00 - pointing check
  • UT 08:02 - pointing check after the infamous co-pointing error on authorization
  • UT 08:09 - initial collimation preset
  • 1" on DX, SX slow to collimate, about 1.1"
  • Sending preset to field - collimation script - no exposures taken
    • script OK - waiting for collimation to get to a good rms
  • UT 08:17 - sending track preset with locked parallactic angle
    • LUCI-1 1052.32 959.9 ---> 924,832
    • LUCI-2 RTD w/Phot Overlay not working - by eye coords: 1105 1003 ---> 977,875
    • both stars visible, changing this to 128x128
    • making LUCI-1 subwin SW 1 988 896 128 128
    • making LUCI-2 subwin SW 1 1041 939 128 128
    • taking another test sub-frame
    • not quite centered
    • still offset - is this because we are rotating fast and center of field isn't center of detector?
    • taking inital exposure again w/HeI and J
    • L1 1039,936 try 64x64 --->1007,904
    • L1 1081 993 try 64x64 --->1049,961
    • LUCI-1 lost collimation badly
  • UT 08:36 - starting - LUCI-1 will be out of collmation
  • Maybe I misunderstood and both stars SHOULD be in it.... Let's take this from the top
  • UT 08:38 - re-sending collimation script
  • UT 08:41 - sending track preset again - re-testing my reading comprehension
    • L1 1020,925----> 892 797
    • L2 1069,1007 (RTD overlay phot worked after a reset)--->941 879
  • UT 08:45 - continuing to next step in script
  • UT 08:46 - starting main part of script - this time everyone was centered up OK (not sure why wasn't last time)
  • NOTE: the luci GUI shows the 5XXX as the last filename read - minor point - perhaps this could be ignored and not passed to the GUI?
  • UT 09:06 - completed
Bright Star Test
  • Negative Dec appeared correctly in the LUCI GUI
  • UT 09:07 - sending preset to TYC 450
    • set subwin Y=1024, X=512 or 1024
  • DX missed the guidestar ---> pointing check
  • UT 09:13 - sending preset again
  • He+z might have been too much - not sure I see the main star
    • Trying FeII + H --> see star
    • L1 1032,1012 -->570,500
    • L2 1029,1051 -->517,539
  • UT 09:24 - starting script looking for crosstalk
  • Do see crosstalk bottom right of star - difference is 108 pixels not a multiple of 64 - if it were from star in view it would be in -line, matches position of HD 289882 1.6' below
  • L1_Crosstalk.jpgL2_Crosstalk.jpg
  • 100x1.25sec CUBE takes ~2min45sec from start to save complete
  • Cross-talk has not dropped off - remains constant from when sub-framing started to current exposure
  • UT 09:37 - completed

Arp 187 - Science LUCI

  • Seeing has improved and stayed good through LUCI tests (<1")
  • UT 02:39 - sending preset
  • LUCI-2 MOS error
    • no mask in FPU
    • Rotator 340
    • Elevation
    • storage to turnout, translator to mask , postcondition violation, motor at required position
    • grabbing 0.5" mask Slot 7 --> position should be -4748, we are at -4748, should be at 275.5, we are at 275.01, tolerance is 0.5
    • putting us right side up to 210
    • initialized MOS --> OK we know where we are, position -129, angle 310.58
    • Manually putting 0.5" slit mask in FPU
  • UT 09:52 - resending preset
  • NUTS! forgot to turn subwindowing off
    • putting mask in turn out, doing this manaully
    • seeing blew up to 2"!!!! AAAHHHH!!!!
    • will take telluric and monitor guiders
    • L1 -9.77", 1.23"
    • L2 -9.166". -1.63"
    • seeing settled back down to <1"
  • UT 10:06 - starting spectra
    • L1 0058-0059
    • L2 0046-0047
  • UT 10:13 - acquiring target
    • weird...why is thru-slit J-band? but field image is H-band...don't recall this last time I ran script with V2
    • L1 -0.3746", 1.4218"
    • L2 0.4094", -1.2947"
  • UT 10:23 - starting spectra
  • weird...it didn't start the grating setup ---> did i click go to fast?
  • UT 10:34 - guiders show 0.8"-0.9"
  • UT 10:46 - guiders bouncing between 0.6"-1"
  • object has Pa-alpha in rotation, but no Br Gamma (it shouldn't, its a dead galaxy) - also shows [SIII] in rotation...outflows?
  • UT 11:02 - seeing starting to get wonky but finishing up now
    • L1 0064-0075
    • L2 0052-0063
  • UT 11;03 - putting LUCIs in safe config

LBC Timing Tests - GPS satellites w/TMS

  • 11:04 - reconfiguring for LBCs
  • 11:10 - Andrew is turning on the TMS lasers and I copied the reference from last night into /tmp
[telescope@robs 20211117]$ cp telescope_left_tmsoffset.dat  /tmp
[telescope@robs 20211117]$ cp telescope_right_tmsoffset.dat  /tmp
[telescope@robs 20211117]$ cp tms_sx_ref.dat /tmp
[telescope@robs 20211117]$ cp tms_dx_ref.dat /tmp
  • 11:11 checks on interval between clicking "PLAY" to resume a paused script and the start of the exposure: 21-22 s blue and 16-17 s red
tail -f lbc.log | grep -e "R CAMERA   CCDCTRL  integration \.\.\." -e "received >PLAY 43 1.00< command from 192.168.34.22 address"
or
tail -f lbc.log | grep -e "B CAMERA CCDCTRL integration \.\.\." -e "received >PLAY 23 1.00< command from 192.168.34.22 address"

  • 11:34 Having run TMS to use the old reference model, paused TMS and ran FPIA to update focus
 -2948   -173   -640   -136   -217    -76     15   1.08   1.19    blue
 -2744     20    237    156    -73    151      3   0.79   1.09     red
    • FPIA is needing 2 iterations on SX to remove Z6. There was some Z5, Z6 on SX "total collimation" when we started - that is unusual

  • 11:41 Removed instrument offsets and also sent the LBC in-focus OB. But I started the TMS loop without first setting a new reference. We need to redo FPIA
  • 11:45 FPIA converged in 1 iteration.
  • 11:46 Removed instrument offsets and set a new reference: ./lbc_set_tms_offset.py -s 2
  • Play OB for 11:50 lbcr.2021118.114950.fits has satellite

  • 11:52 Play script on G05 for 11:55

  • Before we had not synched rsci (updated sync) with gps0 time server. At 12:02 I did. The folowing are images which caught G05 on chip 2:
    • 121505 chip2
    • 121956 chip2
    • 122503 chip2

  • Now I synched bsci with gps0 (NOTE: blue is set to update every 10 min but red is set to update every 1 week)
  • And we confirm that CMU is close to gps0:
lbccontrol@cmu ~]$ ntpdate -q gps0
server 192.168.18.20, stratum 1, offset -0.000359, delay 0.03029
18 Nov 12:29:11 ntpdate[31940]: adjust time server 192.168.18.20 offset -0.000359 sec

  • 12:36 The SX shell ripped - the object is low, below 40 deg, but not below 30 deg. Josh is moving to elevation 42 to reset it. Christian is preparing scripts for R03.
  • 12:41 Slewing to area of new field --- R03 (note that object name is sitll G05)
    • 125455 on chip2
    • 125958 on chip2
    • 130501 on chip2
    • 131001 on chip2
    • 132003 on chip2

Getting too light

Josh is closing up.


-- BarryRothberg - 17 Nov 2021
Topic revision: r5 - 18 Nov 2021, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback